The density of a gas is 1.964 g dm-1 at 273 k and 76 cm Hg. The gas is
(A) CH4
(B) C2H6
(C) CO2
(D) Xe

Grade:12

1 Answers

Adarsh kumar
6 years ago
Density=PM/RT==> M=dRT/P==>M=(1.964×10^(-3) × 82.1×273)/1==>M=44gTherefore, the gas is co2.76cmHg is approximately equal to 1 bar. Therefore, pressure taken is 1 bar.